OpenStudy (tainted):

Medal will be given! How many grams of water vapor (H2O) are in a 10.2 liter sample at 0.98 atmospheres and 26°C? Show all work used to solve this problem

OpenStudy (anonymous):

I'm trying to figure it out.. I just did stuff like this in my Chemistry class

OpenStudy (anonymous):

i have no idea O.o i don't have chemistry...

OpenStudy (anonymous):

Oakay, well what I know is all temperatures should be in Kelvin not Celsius..

OpenStudy (anonymous):

I'm trying to remember if you add 240 or 270

OpenStudy (anonymous):

Wow why am I having trouble with this, I'm struggling i'm not sure how to do this, I can't remember

OpenStudy (anonymous):

can you tell me what you're trying to find?

OpenStudy (tainted):

n=pV/RT 1 mole of h2O = 18g 18*n= x grams

OpenStudy (anonymous):

oh you're doing mole stuff yucky, hold up I'm going thru my notes

OpenStudy (anonymous):

I'm sorry I do not know..

OpenStudy (jfraser):

To convert C into K, add 273. Plug into the ideal gas law and solve for moles of water. convert moles of water into grams of water

OpenStudy (tainted):

Like so? m = PVM / RT R = (1atm x 22.4L) / (1mole x 273K) = 0.0821Latm/moleK T = 26 + 273 = 299K m = 0.98atm x 10.2L x 18g/mole / (0.0821Latm/molek x 299K) = 7.33gH2O

OpenStudy (jfraser):

i get about 7.5g, but that's close enough, given differences in rounding

OpenStudy (tainted):

Ok thanks!

OpenStudy (jfraser):

yw